OpenStudy (lacris):

Please help me :) I will give you a shiny medal :3 A probability experiment is conducted in which the sample space of the experiment is S={2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13}, Event E = {2, 3, 4, 5, 6, 7} and event G= {9, 10, 11, 12} Assume that each outcome is equally likely. List the outcomes in E and G. Are E and G mutually exclusive? E and G = {?} I placed:

OpenStudy (lacris):

First I did: E and G {2, 3, 4, 5, 6, 7, 9, 10, 11, 12} But I got that incorrect... Isn't it mutually exclusive? (Or disjointed?) such as P(E or G) = P(E) + P(G)?

OpenStudy (lacris):

@agent0smith I hope your not busy :>

OpenStudy (agent0smith):

Math doesn't read easily in paragraph form: S={2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13} E = {2, 3, 4, 5, 6, 7} G= {9, 10, 11, 12} E and G = { } the empty set and thus they are mutually exclusive (there's zero overlap)

OpenStudy (agent0smith):

E and G means only the elements that show up in BOTH E and G, of which there are none. What you did first is E or G, which is all of the elements in E and G, combined together.

OpenStudy (agent0smith):

P(E or G) = P(E) + P(G) Yes this is true for mutually exclusive, since you'd normally use P(E or G) = P(E) + P(G) - P(E and G) but when two events are mutually exclusive, P(E and G) = 0, so P(E or G) = P(E) + P(G) - 0

OpenStudy (lacris):

@agent0smith thank you so much for helping me understand! Sorry for the late reply, but it means a lot that you took time to help!